OpenStudy (anonymous):

Ashe went from one hundred sixty pounds to one hundred eighty pounds once he started a weight-training program at his gym. What was Ashe's percent increase in weight? eight percent eleven point one percent twelve point five percent twenty percent

OpenStudy (anonymous):

I'll be back I know the answer

OpenStudy (anonymous):

@radar

OpenStudy (calculusxy):

\[\frac{ current - starting } { starting } \times 100\]

OpenStudy (texaschic101):

(new amount - original amount) / original amount ...then multiply by 100 (180 - 160) / 160 = 20/160 = 0.125 0.125 * 100 = 12.5%

OpenStudy (calculusxy):

do you understand now?

OpenStudy (anonymous):

half way

OpenStudy (calculusxy):

what is confusing for you?

OpenStudy (anonymous):

put it into percent

OpenStudy (calculusxy):

well the current value is the new amount (which in this case is 180lbs). the starting value (which was the original value - is 160lbs). so you subtract those two first and get the difference of 20. then you divide 20 by 160 (the starting value) and get 0.125 . when you multiply that by 100, you get 12.5% .

OpenStudy (calculusxy):

when changing a decimal (0.125) to percents, you have to multiply by 100 since percents are out of 100. an easy way to multiply by 100, is to just move the decimal place two times to the right. therefore, when moving the decimal two places to the right, we have 12.5%

OpenStudy (calculusxy):

do you now understand the procedure?

OpenStudy (anonymous):

yes an thank u

OpenStudy (calculusxy):

no problem :)